OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=4/x and g(x)=4/x

OpenStudy (anonymous):

@jim_thompson5910 this is the last question on my homework for the weekend... Please help me!!!

jimthompson5910 (jim_thompson5910):

How far did you get?

OpenStudy (anonymous):

I have no clue how to start this one... It says they're supposed to be inverses but they''re the same...

jimthompson5910 (jim_thompson5910):

what is f(g(x))?

OpenStudy (anonymous):

x

jimthompson5910 (jim_thompson5910):

well that's what you have to show

jimthompson5910 (jim_thompson5910):

but how would you show it?

OpenStudy (anonymous):

by multiplying f(x) and g(x)?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

that's if they wanted f(x) * g(x)

OpenStudy (anonymous):

okay then I'm not entirely sure...

jimthompson5910 (jim_thompson5910):

an example of f(g(x)) let f(x) = 2x+3 and g(x) = x^4 --------------------------------- f(x) = 2x + 3 ... start with the outer function f(g(x)) = 2*g(x) + 3 ... replace ALL copies of x with g(x) f(g(x)) = 2x^4 + 3 ... replace ALL copies of g(x) on the right side with x^4 (since g(x) = x^4)

jimthompson5910 (jim_thompson5910):

So that means if f(x) = 2x+3 and g(x) = x^4, then f(g(x)) = 2x^4 + 3

OpenStudy (anonymous):

okay, I got \[f(g(x))=\frac{ 4 }{ \frac{ 4 }{ x } }\] but I think that's wrong...

OpenStudy (the_fizicx99):

Lol I have a doubt, I got 4/4/x which I'm not sure is right. But knowing that they're asking they they both equal the same g(f(x)) would also = 4/4/x which make me unsure of my answer. I'm 50/50 :p

OpenStudy (anonymous):

yeah, that's the same thing I got.... @jim_thompson5910 is that right?

jimthompson5910 (jim_thompson5910):

think of that first uppermost 4 as 4/1

jimthompson5910 (jim_thompson5910):

you will then have 4/1 over 4/x so divide those fractions and tell me what you get

OpenStudy (anonymous):

okay...?

OpenStudy (anonymous):

that's the same as 4/1 * x/4 right?

OpenStudy (the_fizicx99):

Well i think its actually right, as it would be 4 would be 4/1 divided by 4/x or 4/1 Or 4/1 * 1/4 = 4/4 = 1 Idk lol I was learning the inverse at the same time which is what set me off :P

OpenStudy (anonymous):

OH! The 4s cancel out and you're left with 1x or just x!!!! right?

jimthompson5910 (jim_thompson5910):

keep going Sunshine447

jimthompson5910 (jim_thompson5910):

good, you got it

OpenStudy (the_fizicx99):

Yes :)

OpenStudy (the_fizicx99):

@jim_thompson5910 , how do you say f(^-1) ? I keep forgetting it.

OpenStudy (anonymous):

now how does this make them inverse?

jimthompson5910 (jim_thompson5910):

you've just shown that f(g(x)) = x now you need to show that g(f(x)) = x

jimthompson5910 (jim_thompson5910):

if you can show that f(g(x)) = x and g(f(x)) = x are both true, then f(x) and g(x) are inverses of one another

OpenStudy (anonymous):

ut's the same thing right? the 4/1*x/4 and canceling the 4s and being left with x?

jimthompson5910 (jim_thompson5910):

pretty much since f(x) = g(x)

jimthompson5910 (jim_thompson5910):

tHe_FiZiCx99 I typed in f^{-1}(x)

OpenStudy (the_fizicx99):

No, I meant how t\do you verbally say it ~ f of inverse x ~ ?

jimthompson5910 (jim_thompson5910):

oh you say "f inverse"

OpenStudy (the_fizicx99):

Ohh, ty :>

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Thank you sosososososososososo much!!!!

jimthompson5910 (jim_thompson5910):

sure thing